# Dzyalonshinskii-Moriya interaction in Fe5GeTe2 epitaxial thin films

### v1

@@ Abstract
Van der Waals ferromagnets, such as Fe5GeTe2, offer a promising platform for spintronic devices based on chiral magnetic textures, provided a significant Dzyaloshinskii-Moriya interaction (DMI) can be induced to stabilise them. Here, we directly measure DMI in epitaxial Fe5GeTe2 thin films using Brillouin light scattering spectroscopy and observe a consistent DMI ($D=0.04$ mJ/m2) across various thicknesses. Its weak thickness dependence, combined with the nominally symmetric film interfaces, suggests a bulk origin. Although we do not determine the microscopic mechanism, our findings are compatible with ab-initio calculations linking DMI to partial ordering of Fe split sites. Additionally, we find a low magnetic dissipation ($\alpha<0.003$). The observed DMI, which could be further enhanced by optimising the Fe site ordering, combined with low dissipation, makes Fe5GeTe2 a strong candidate for exploring the dynamics of chiral magnetic textures in two-dimensional materials.
@@ 1 Introduction
Moreover, there have been reports of chiral magnetic structures in FexGeTe2, opening the way to their use in the many proposed devices based on chiral textures such as skyrmions. Homo-chiral spin spirals were observed in exfoliated flakes of Fe3GeTe2 2 (thickness $t=185$ nm), and Néel skyrmions were reported in Fe3GeTe2 (70 nm)/[Co/Pd] 3 and WTe2/Fe3GeTe2 (32 nm) 4 heterostructures. In other members of the FexGeTe2 family, Néel skyrmions were observed in Fe5-xGeTe2 5, 6 (on 61 and 100 nm thick flakes), as well as in (FeCo)5GeTe2 (110 nm) 7. On the other hand, other groups working also on flakes of Fe5GeTe2 observed only non-chiral Bloch bubbles 8, 9.
@@
The observation of homo-chiral textures suggests a significant Dzyaloshinskii-Moriya interaction (DMI). Indeed, a large DMI parameter ($D=1.0$ mJ/m2) was estimated in ref. 4. A nonzero DMI requires a symmetry-breaking mechanism along the out-of-plane direction. However, both Fe3GeTe2 and Fe5GeTe2 are presumed to have inversion-symmetric crystal structures incompatible with DMI10. Nevertheless, several experimental studies suggest that the actual crystalline structure of FexGeTe2 may lack inversion symmetry. Possible sources of symmetry breaking include Fe split-site ordering 11, Fe self-intercalation in the van der Waals gaps 12, ordered Te vacancies 9, and asymmetric Fe vacancy distribution 13. Additionally, deviations from the expected vdW stacking sequence (ABCABC in Fe5GeTe2 and ABAB in Fe3GeTe2) could also break inversion symmetry.
@@
Additionally, large DMI has been observed in ultra-thin magnets with a symmetric bulk structure, such as nanometric Co films in a AlOx/Co/Pt triplet film (up to $D=2$ mJ/m2 14). In these cases, the DMI does not arise from the bulk crystalline structure but instead from asymmetric interfaces of the ultra-thin magnetic layer, which can be engineered to increase the effect. Although this interfacial mechanism could be expected to be weaker in vdW materials due to their weak interlayer coupling, first-principle calculations by Park et al. 15 show that the interface of Fe3GeTe2 with its natural oxide can induce a significant DMI ($\approx 2$ mJ/m2 for a monolayer of Fe3GeTe2). Interfacial mechanisms, however, are only significant in ultra-thin systems and all the cited experimental observations were done in relatively thick (30 to 200 nm) exfoliated flakes.
@@
Characterizing the DMI solely from the observation of magnetic textures is challenging, as these are influenced by multiple interactions, including dipolar effects, magnetocrystalline anisotropy, and exchange. In some cases, dipolar interactions may even create seemingly chiral textures at the surface of thick samples 3, 16. To determine the origin of chiral magnetic textures and disentangle interfacial and bulk contributions, a more direct measurement of the DMI is essential. In this letter, we investigate the static and dynamic magnetic properties of epitaxial Fe5GeTe2 thin films of varying thickness using Brillouin light scattering (BLS) spectroscopy. By directly measuring the DMI, we unambiguously demonstrate the presence of a significant DMI of bulk origin. Furthermore, we show that the magnetic damping in these films is low compared to conventional metallic ferromagnets. These properties make Fe5GeTe2 a promising material for the study of the propagation of chiral magnetic textures.
@@ 2.1 Sample growth & structure
Fe5GeTe2 films ($t=$ 12, 8 and 3 nm) were epitaxially grown on Al2O3(0001). The 12-nm-thick film studied here was characterised in depth by Ribeiro et al. 17, and all films were grown following the protocol reported in that work. The films were capped by 3 nm of Al deposited at room temperature, then naturally oxidized in air. This capping layer prevents the oxidation of Fe5GeTe2 and provides symmetric top and bottom interfaces (Fig. 1a). The good crystalline quality of the 12-nm film was observed by X-ray diffraction (XRD) and scanning transmission electron microscopy 17. In line with this result, the XRD scans in Fig. 1b demonstrate a well-defined Fe5GeTe2 lattice for all layers. From these XRD measurements, the thickness of one Fe5GeTe2 monolayer is found equal to 0.98 nm, independently of the film thickness. The RHEED patterns of Fe5GeTe2 (8 nm) in the (100) and (110) reciprocal directions, measured after growth, are shown in Fig. 1c. Similar anisotropic diffraction patterns are observed at all thicknesses and demonstrate the good epitaxy on sapphire. However, faint (10) diffraction rods are visible in the (110) direction (peak at $q/q(11)=1/\sqrt{3}$ in Fig. 1d). They denote the presence of a small proportion of domains that are rotated by 30° from the main crystal orientation. This feature is more apparent for the thinnest layer and disappears for the 12 nm film, indicating an improvement of the crystal in-plane orientation quality with increasing thickness. In addition, the RHEED patterns display a ($3\times 1$) reconstruction characteristic of a ($\sqrt{3}\times\sqrt{3}$) R30° superstructure (peaks at q/q(11) = 1/3 and 2/3 in Fig. 1d). This superstructure is observed in all three films, and is more apparent at larger thicknesses.
Fe5GeTe2 is expected to have two partially-filled Fe sites (“split sites”) at different vertical positions in the elementary monolayer, "U" or "D" (see Fig. 1a). The $\sqrt{3}\times\sqrt{3}$ superstructure may be attributed to a large-scale lateral ordering of the Fe split sites in a “UUD” or “UDD” pattern (inset of Fig. 1d), which was previously experimentally observed by Ly et al. 11 and found to be more energetically stable than the “UUU” ordering by ab initio calculations 18. This superstructure is also compatible with an ordering of Te vacancies as reported by Gopi et al. 9. Both these structures break the inversion symmetry of the crystal. The presence of this superstructure in the deeper layers of the films (beyond their top surface) cannot be assessed from these RHEED measurements, which were recorded after the deposition.
@@ 2.2 Magnetic properties
The spontaneous magnetization ($M_{s}$) of the three samples was measured by SQUID magnetometry as a function of temperature (Fig. 2). Thinner films showed lower $M_{s}$, with the 3-nm film showing about one-third of the $M_{s}$ of the 12-nm film at 150 K (see Table 1). The Curie temperature ($T_{C}$) of the thicker films (297–299 K) is similar to bulk values (270–310 K 19), whereas the 3-nm film had a lower $T_{C}$ (258 K). The evolution of $M_{s}$ and $T_{c}$ indicates that a thickness of a few layers is needed to fully develop the bulk magnetic properties of the films. The $M_{s}(T)$ curves showed a monotonic trend, unlike previous reports of transitions in $M_{s}(T)$ 1, 8, 11, 20. However, those studies focused on virgin films, whereas our samples had been previously magnetized. These transitions may also be specific to bulk samples rather than epitaxially-grown films.
Hysteresis loops measured at 150 K by longitudinal magneto-optical Kerr effect (LMOKE) (Fig. 2b) show in-plane magnetization reversal with a small coercive field of a few mT around 150 K. The sharp transitions in the loops indicate reversal by nucleation and domain wall propagation, typical of homogeneous films. LMOKE imaging of the 12-nm film (Fig. 2c) confirms domain wall propagation during reversal and reveals large domains (10–100 $\mu$m), further indicating film homogeneity.
@@
The g-factor obtained from the BLS measurements (see Methods) consistently falls in the range $g=2.1-2.3$, showing no significant temperature dependence, in close agreement with previous reports 21, 1. The notable deviation from the free-electron value of $g=2.0$ suggests a substantial orbital moment and significant spin-orbit coupling in the material. This is consistent with the orbital/spin moment ratio measured in 17.
The Gilbert dissipation parameter ($\alpha$) can be inferred from the BLS peak width ($\Gamma$), similar to its extraction from ferromagnetic resonance measurements. However, $\Gamma$ is also broadened by inhomogeneities in the film’s magnetic properties, which are typically separated using high magnetic fields due to their distinct field dependence. Since the available field was insufficient for this separation, we estimated an upper bound, $\alpha_{\rm eff}$ (see Fig. 5 and Methods). The lowest $\alpha_{\rm eff}$ values for each film are listed in Table 1. The $\alpha_{\rm eff}$ is larger for thinner films, which could be due to a higher dissipation, as is often observed in other magnetic films (e.g. 22). It may also be due to a higher inhomogeneity of the magnetic properties, as the properties of thinner films are more affected by variations of the thickness or of the interfaces. Overall, $\alpha_{\rm eff}$ as low as $3\cdot 10^{-3}$ were observed, which is lower than most typical metals ($\sim 10^{-2}$) and lower than reported measurements in bulk Fe5GeTe2 1. A low magnetic dissipation parameter enables the motion of textures with lower applied torques, such as skyrmions driven by spin currents, adding to the interest of Fe5GeTe2 as a material to study chiral textures.
Figure 3: a) Example of a BLS spectrum of the 12 nm film, in black, taken at $\phi=45$º (150 K, $B=130$ mT). The green lines show the fit to a Lorentzian peak (with the fitted centres written above the peaks). The central peak corresponds to non-scattered light and shows the instrumental width. b) Schematic of the BLS geometry. c) Stokes/anti-Stokes frequency difference, $\Delta f$, versus wave vector $k$ for the three films (points; taken at $B=$ 211, 250, and 130 mT, respectively.) The linear fits (lines) are used to extract $v_{\rm DMI}$ and $D$. The offset of the fit accounts for a possible imperfect alignment of the spectrometer. d) Extracted $D$ versus sample thickness $t$ (points) and fits assuming a uniform bulk origin (constant blue line) and an interfacial origin ($\propto 1/t$, green line).
2.3 DMI
The Stokes and anti-Stokes peaks in the BLS spectrum (Fig. 3a) correspond to cycloidal spinwaves of opposite rotation sense (Fig. 3b), and are subject to an opposite energy contributions from the DMI. In the limit of thin films ($kt\ll 1$, where $k$ is the wave vector of the spinwave determined by the incidence angle of the light), as is the case here, the difference in frequency between the Stokes and anti-Stokes peaks ($\Delta f\equiv|f_{\rm anti-Stokes}|-|f_{\rm Stokes}|$) can be fully attributed to this effect, and can be used to determine the DMI parameter $D$ 23, 14:
$\Delta f=\frac{2\gamma}{M_{s}}Dk,$ (1)
where $\gamma=g\mu_{B}/\hbar$ is the gyromagnetic ratio (and $\mu_{B}$ and $\hbar$ are the Bohr’s magneton and the reduced Planck constant, respectively.) As there may be a small constant $\Delta f$ due to an imperfect alignment of the spectrometer, several spectra were measured and the slope of $\Delta f(k)$ (which we dub $v_{\rm DMI}$) was determined. The expected reversal of the sign of $\Delta f$ with a reversal of the field direction was also verified. Fig. 3c shows the measured $\Delta f(k)$ for the 3, 8 and 12 nm films at 150 K. We observe that there is clearly a linear slope and thus a finite $D$ in all three films. All three show a slope, and thus a DMI, of the same sign. With the geometry of our setup ($\bm{\mathbf{k}}_{\rm Stokes}\times\bm{\mathbf{B}}$ pointing into the film surface, Fig. 3b), this positive slope corresponds to favoured right-handed spinwaves, and thus a positive DMI parameter. The calculated $D$ parameters (using eq. 1) are shown in Tab. 1 for the three films.
As a verification, the DMI was also measured in one of the films ($t=8$ nm) at room temperature, and no DMI was observed, as expected when approaching $T_{C}$.
The measured DMI varies between $D=0.02$ to 0.05 mJ/m2 and is comparable to the values typically observed in spintronic systems, although lower than the highest known effects (for reference, the strong interfacial DMI in Co/Pt would yield approximately 0.2 mJ/m2 in a 8-nm-thick film 14).
3 Discussion
The observation of DMI requires broken inversion symmetry along the $c$-axis, either in the bulk of the material or due to asymmetric interfaces. This symmetry breaking must persist over a large lateral scale, as BLS measurements average over the 30-µm laser spot and are consistent between points several 10s of µm apart. The dependence of $D$ on film thickness $t$ can help distinguish between bulk and interfacial origins: a bulk contribution would result in a constant $D$, whereas an interfacial contribution should scale as $D\propto 1/t$ 14. Figure 3d shows $D(t)$ along with best-fit curves for both scenarios. Clearly, the data do not follow the $1/t$ trend expected for interfacial DMI. This finding aligns with the rather symmetric environment of our films, which have two electronically similar interfaces (Al2O3/Fe5GeTe2 and Fe5GeTe2/AlOx). Instead, the $D(t)$ data is more compatible with the constant behaviour of a bulk origin, with a $D=0.04$ to 0.05 mJ/m2, as measured in the thickest films. The weaker DMI in the 3-nm-thick film, which correlates with lower $M_{s}$ and $T_{c}$, can be attributed to weaker ferromagnetism in the ultrathin regime.
[…]
@@ 3 Discussion
A second mechanism involves symmetry breaking within the Fe5GeTe2 vdW layers. This is supported by our RHEED observations of a $\sqrt{3}\times\sqrt{3}$ superstructure, which cannot derive from a specific vdW stacking. A non-centrosymmetric phase with a $\sqrt{3}\times\sqrt{3}$ superstructure was reported by Gopi et al. 9, who ascribed it to a pattern of 1/3 Te vacancies ordered along the perpendicular direction. Since this phase represented only a small fraction ($<10^{-2}$) of the studied bulk crystals, it was deemed to have a negligible effect on the material’s DMI. While this phase might be more prevalent in epitaxial films, Rutherford backscattering analysis of our films 17 showed no significant Te depletion, suggesting that this mechanism is unlikely to explain our results 222 The Te stoichiometry in the 12-nm film was measured to be Te2±0.08 17. A full pattern of 1/3 Te vacancies in one of the sides of all vdW layers (i.e., a loss of 1/6 of the expected Te) would give Te1.7..
@@
Other groups interpreted this $\sqrt{3}\times\sqrt{3}$ superstructure as the ordering of Fe split sites (“D” and “U” sites in Fig.1a,d) 11, 19. In the centrosymmetric structure, these sites are equally and randomly populated. However, ab initio calculations showed that the “UDD” and “UUD” configurations are energetically more favorable 18. While our RHEED measurements cannot determine whether this ordering extends beyond the top layer, electron microscopy analyses in Ref. 19 showed a similar short-range order in the bulk of Fe5GeTe2. Three groups have calculated the DMI assuming a perfectly ordered “UUU” arrangement in a monolayer, using different computational methods 26, 27, 5. Their predicted $|D|$ values vary significantly (0.76, 0.6, and 0.06 mJ/m2), despite all employing commonly used techniques. Our measured $|D|$ (0.04 – 0.05 mJ/m2) is comparable to the lowest theoretical estimate, and much lower than the highest prediction. This quantitative difference may be due to several reasons. First, the calculations do not account for temperature, which reduces the macroscopic $D$. Second, our films are not ordered in the “UUU” configuration since this would be inconsistent with the observed $\sqrt{3}\times\sqrt{3}$ superstructure. According to Ghosh et al. 27, the DMI of the “UUD” or “UDD” configurations is only slightly lower than that of the “UUU” configuration ($|D|=0.05$ instead of 0.06 mJ/m2), contrary to the naïve expectation of a 2/3 reduction. Finally, by symmetry, the “UUD” and “UDD” configurations contribute equally but with opposite sign DMI. While one of the two must be dominant to account for the observed DMI, the other may be present in a smaller amount and partially compensate the net DMI.
@@
Assuming that the DMI arises from an ordering of the Fe split sites, the sign of $D$ can be used to infer the dominant split-site configuration. Both refs. 26 and 5 found that a fully-ordered “UUU” configuration favours counterclockwise spinwaves, corresponding to a negative DMI. Therefore, the observed positive DMI suggests that the predominant ordering in our films is “UDD.”
The presence of a symmetry-breaking mechanism extending both laterally over large scales and through the film thickness, combined with the consistent DMI sign observed in all three films, indicates a robust and reproducible effect. This suggests that the symmetry breaking is likely associated with the growth process, which is inherently asymmetric.

### v2

# Dzyaloshinskii-Moriya interaction in Fe5GeTe2 epitaxial thin films

@@ Abstract
Van der Waals ferromagnets, such as Fe5GeTe2, offer a promising platform for spintronic devices based on chiral magnetic textures, provided a significant Dzyaloshinskii-Moriya interaction (DMI) can be induced to stabilise the textures. Here, we directly measure DMI in epitaxial Fe5GeTe2 thin films using Brillouin light scattering spectroscopy and observe a consistent DMI ($D=0.04$ mJ/m2) across various thicknesses. Its weak thickness dependence, combined with the nominally symmetric film interfaces, suggests a bulk origin. Although we do not determine the microscopic mechanism, our findings are compatible with ab initio calculations linking DMI to partial ordering of Fe split sites. Additionally, we find a low magnetic dissipation ($\alpha<0.02$). The observed DMI, which could be further enhanced by optimising the Fe site ordering, combined with low dissipation, makes Fe5GeTe2 a strong candidate for exploring the dynamics of chiral magnetic textures in two-dimensional materials.
@@
Moreover, there have been reports of chiral magnetic structures in FexGeTe2, opening the way to their use in the many proposed devices based on chiral textures such as skyrmions. Homo-chiral spin spirals were observed in exfoliated flakes of Fe3GeTe2 2 (thickness $t=185$ nm), and Néel skyrmions were reported in Fe3GeTe2 (70 nm)/[Co/Pd] 3 and WTe2/Fe3GeTe2 (32 nm) 4 heterostructures. In other members of the FexGeTe2 family, Néel skyrmions were observed in Fe5-xGeTe2 5, 6 (on 61 and 100 nm thick flakes), as well as in (Fe,Co)5GeTe2 (110 nm) 7 or in the related Fe3-xGaTe2 8, 9. On the other hand, other groups working also on flakes of Fe5GeTe2 observed only non-chiral Bloch bubbles 10, 11.
@@
The observation of homo-chiral textures suggests a significant Dzyaloshinskii-Moriya interaction (DMI). Indeed, a large DMI parameter ($D=1.0$ mJ/m2) was estimated in ref. 4. A nonzero DMI requires a symmetry-breaking mechanism along the out-of-plane direction. However, both Fe3GeTe2 and Fe5GeTe2 are presumed to have inversion-symmetric crystal structures incompatible with DMI12. Nevertheless, several experimental studies suggest that the actual crystalline structure of FexGeTe2 may lack inversion symmetry. Possible sources of symmetry breaking include Fe split-site ordering 13, Fe self-intercalation in the van der Waals gaps 14, ordered Te vacancies 11, and asymmetric Fe vacancy distribution 15. Additionally, deviations from the expected vdW stacking sequence (ABCABC in Fe5GeTe2 and ABAB in Fe3GeTe2) could also break inversion symmetry.
@@
DMI can also arise from asymmetric interfaces in thin magnetic layers. This mechanism has been well established in metallic systems, such as nanometric Co films in a AlOx/Co/Pt triplet film (up to $D=2$ mJ/m2 16, 17). Although this interfacial mechanism could be expected to be weaker in vdW materials due to their weak interlayer coupling, first-principle calculations by Park et al. 18 show that the interface of Fe3GeTe2 with its natural oxide can induce a significant DMI ($\approx 2$ mJ/m2 for a monolayer of Fe3GeTe2). Interfacial mechanisms, however, are only significant in ultra-thin systems, since their effects are progressively diluted with increasing thickness, while all the cited experimental observations involve relatively thick (30-200 nm) exfoliated flakes.
@@
Characterizing the DMI solely from the observation of magnetic textures is challenging, as these are influenced by multiple interactions, including dipolar effects, magnetocrystalline anisotropy, and exchange. In some cases, dipolar interactions may even create seemingly chiral textures at the surface of thick samples 3, 19. To determine the origin of chiral magnetic textures and disentangle interfacial and bulk contributions, a more direct measurement of the DMI is essential. In this letter, we investigate the static and dynamic magnetic properties of epitaxial Fe5GeTe2 thin films of varying thickness using Brillouin light scattering (BLS) spectroscopy. By directly measuring the DMI, we unambiguously demonstrate the presence of a significant DMI of bulk origin. Furthermore, we show that the magnetic damping in these films is low for a metallic ferromagnet. These properties make Fe5GeTe2 a promising material for the study of the propagation of chiral magnetic textures.
@@ 1.1 Sample growth & structure
Fe5GeTe2 films ($t=$ 12, 8 and 3 nm) were epitaxially grown on Al2O3(0001) at 350°C by coevaporation of elemental Fe, Ge and Te, with a Fe:Ge flux ratio of 5 and with a slightly overstoichiometric Te flux, and then annealed $\sim 500$°C for 10 min. The 12-nm-thick film was characterised in depth in Ribeiro et al. 20, and all films were grown following the protocol reported in that work. The films were capped by 3 nm of Al deposited at room temperature, then naturally oxidized in air. This capping layer prevents the oxidation of Fe5GeTe2 and provides similar top and bottom interfaces (Fig. 1a).
The X-ray diffraction (XRD) specular scans (Fig.1b) show single-phase Fe5GeTe2 with a constant $c$-axis lattice parameter (0.98 nm per monolayer) across all thicknesses (see S.I.), consistent with the good crystalline quality previously observed for the 12-nm film by XRD and scanning transmission electron microscopy (STEM) 20. The RHEED patterns of Fe5GeTe2 (8 nm) in the (100) and (110) reciprocal directions, measured after growth, are shown in Fig. 1c. Similar anisotropic diffraction patterns are observed at all thicknesses and demonstrate the good epitaxy on sapphire. However, faint (10) diffraction rods are visible in the (110) direction (peak at $q/q(11)=1/\sqrt{3}$ in Fig. 1d). They denote the presence of a small proportion of domains that are rotated by 30° from the main crystal orientation. This feature is more apparent for the thinnest layer and disappears for the 12 nm film, indicating an improvement of the crystal in-plane orientation quality with increasing thickness.
In addition, the RHEED patterns display a ($3\times 1$) reconstruction characteristic of a ($\sqrt{3}\times\sqrt{3}$) R30° superstructure (peaks at q/q(11) = 1/3 and 2/3 in Fig. 1d). This superstructure is observed in all three films, and is more apparent at larger thicknesses. The nature of this feature will be discussed later.
@@ 1.2 Magnetic properties
The spontaneous magnetization ($M_{s}$) was measured by SQUID magnetometry versus temperature (Fig. 2). Thinner films showed lower $M_{s}$, with the 3-nm film showing about one-third of the $M_{s}$ of the 12-nm film at 150 K (see Table 1). The Curie temperature ($T_{C}$) of the thicker films (297–299 K) is similar to bulk values (270–310 K 1), whereas the 3-nm film had a lower $T_{C}$ (258 K). The evolution of $M_{s}$ and $T_{c}$ indicates that a thickness of a few layers is needed to fully develop the bulk magnetic properties of the films. The $M_{s}(T)$ curves showed a monotonic trend, unlike previous reports of various transitions in $M_{s}(T)$ 21, 10, 13, 22. However, those studies focused on virgin films, whereas our samples had been previously magnetized. These transitions may also be specific to bulk samples rather than epitaxially-grown films, suggesting structural or chemical differences between these two classes of materials.
[…]
@@
The samples were further characterised by studying the spin wave frequency as a function of field and wave vector using BLS spectroscopy (analysed using the laws of dispersion determined in refs. 23, 16, 24; see S.I.). The negative anisotropy field $B_{K}$ extracted from the BLS measurements of frequency versus field (Fig. S2) confirms the easy-plane (i.e., ab plane) anisotropy (at 150 K; Tab. 1). Its magnitude is significantly larger than the expected contribution from shape anisotropy (given by $\mu_{0}M_{s}$), a behavior previously observed in Fe5GeTe2 20, 21. This suggests either the presence of an intrinsic easy-plane magnetocrystalline anisotropy or that the magnetisation is inhomogeneous across the vdW layers and gaps, leading to a stronger shape anisotropy than expected from the measured $M_{s}$ (see S.I.).
The g-factor obtained from the BLS measurements consistently falls in the range g=2.12.3g=2.1-2.3, showing no significant temperature dependence, in close agreement with previous reports 25, 21. The notable deviation from the free-electron value of $g=2.0$ suggests a substantial orbital moment and significant spin-orbit coupling in the material. This is consistent with the orbital/spin moment ratio measured in ref. 20.
The Gilbert dissipation parameter ($\alpha$) can be inferred from the BLS peak width ($\Gamma$), similar to its extraction from ferromagnetic resonance measurements. However, $\Gamma$ is also broadened by inhomogeneities in the film’s magnetic properties, which are typically disentangled from the effect of $\alpha$ with high magnetic fields taking advantage of their distinct field dependence. Since the available field was insufficient for this separation, we estimated an upper bound, $\alpha_{\rm eff}$ (see Fig. S3 and the S.I.). The lowest $\alpha_{\rm eff}$ values for each film are listed in Table 1. The $\alpha_{\rm eff}$ is larger for thinner films, which could be due to a higher dissipation, as is often observed in other magnetic films (e.g. ref. 26). It may also be due to a higher inhomogeneity of the magnetic properties, as the properties of thinner films are more affected by variations of the thickness or of the interfaces. Overall, $\alpha_{\rm eff}$ as low as 0.02 were observed, lower than reported measurements in bulk Fe5GeTe2 21.
This low $\alpha$ is noteworthy, as metallic systems containing heavy elements typically exhibit large damping. This behavior is commonly interpreted using the breathing Fermi surface model 27, which predicts that $\alpha$ scales quadratically with the spin–orbit coupling (SOC) strength, linearly with the density of states (DoS) at the Fermi level, and with the electron relaxation time $\tau$ (itself proportional to the conductivity). The DoS of Fe5GeTe2 is reported to be comparable to that of iron 28. Conversely, its low conductivity suggests a short $\tau$. In addition, its weak magnetocrystalline anisotropy points to a modest SOC. Indeed, Te primarily bonds to the weakly magnetic Fe1 sublattice 20, limiting its contribution to damping. These factors may explain the observed $\alpha$, but they also highlight the need for an ab initio study of damping in these materials. Moreover, such low magnetic dissipation facilitates the motion of spin textures, such as skyrmions, further adding to the interest of Fe5GeTe2 as a material for Spintronics.
Figure 3: a) A BLS spectrum of the 12 nm film ($\phi=45$º, 150 K, $B=130$ mT). The green lines show the Lorentzian fit (fitted centres are written above the peaks). The central peak corresponds to non-scattered light and shows the instrumental width. b) Schematic of the BLS geometry. c) Stokes/anti-Stokes frequency difference, $\Delta f$, versus wave vector $k$ for the three films (points; taken at $B=$ 211, 250, and 130 mT, respectively.) The linear fits (lines) are used to extract $v_{\rm DMI}$ and $D$. d) Extracted $D$ versus sample thickness $t$ (points) and fits assuming a uniform bulk origin (constant blue line) and an interfacial origin ($\propto 1/t$, green line).
1.3 DMI
The Stokes and anti-Stokes peaks in the BLS spectrum (Fig. 3a) correspond to cycloidal spinwaves of opposite rotation sense (Fig. 3b), and are subject to an opposite energy contributions from the DMI. In the limit of thin films ($kt\ll 1$, where $k$ is the wave vector of the spinwave determined by the incidence angle of the light), as is the case here, the difference in frequency between the Stokes and anti-Stokes peaks ($\Delta f\equiv|f_{\rm anti-Stokes}|-|f_{\rm Stokes}|$) can be fully attributed to this effect, and can be used to determine the DMI parameter $D$ 29, 16:
$\Delta f=\frac{2\gamma}{\pi M_{s}}Dk,$ (1)
where $\gamma=g\mu_{B}/\hbar$ is the gyromagnetic ratio (and $\mu_{B}$ and $\hbar$ are the Bohr’s magneton and the reduced Planck constant, respectively.) As there may be a small constant $\Delta f$ due to an imperfect alignment of the spectrometer, several spectra were measured and the slope of $\Delta f(k)$ (which we dub $v_{\rm DMI}$) was determined. The expected reversal of the sign of $\Delta f$ with a reversal of the field direction was also verified. Fig. 3c shows the measured $\Delta f(k)$ for the 3, 8 and 12 nm films at 150 K. We observe that there is clearly a linear slope and thus a finite $D$ in all three films. All three show a slope, and thus a DMI, of the same sign. With the geometry of our setup ($\bm{\mathbf{k}}_{\rm Stokes}\times\bm{\mathbf{B}}$ pointing into the film surface, Fig. 3b), this positive slope corresponds to favoured right-handed spinwaves, and thus a positive DMI parameter. The calculated $D$ parameters (eq. 1) are shown in Tab. 1 for the three films. As a check, the DMI of the 8-nm film was also measured at 100 K and 300 K (Fig. S4 in S.I.), yielding a similar value at 100 K and no detectable DMI at 300 K, as expected near $T_{C}$.
2 Discussion
The measured DMI (0.02 to 0.05 mJ/m2) is comparable to values in many spintronic systems, though lower than the strongest known effects (e.g., $\sim$0.2 mJ/m2 for Co/Pt at 8 nm thickness 16). Micromagnetic simulations30 (see S.I.) show that this DMI suffices to induce chiral magnetic textures if the anisotropy is tuned to favour out-of-plane magnetisation.
The observation of DMI that favours cycloidal textures requires broken inversion symmetry along the $c$-axis, either in the bulk of the material or due to asymmetric interfaces. This symmetry breaking must persist over a large lateral scale, as BLS measurements average over the 30-µm laser spot and are consistent between points several 10s of µm apart. The dependence of $D$ on film thickness $t$ can help distinguish between bulk and interfacial origins: a bulk contribution would result in a constant $D$, whereas an interfacial contribution should scale as $D\propto 1/t$ 16. Figure 3d shows $D(t)$ alongside best-fit curves for both scenarios. Clearly, $D(t)$ deviates from the $1/t$ trend expected for interfacial DMI, in alignment with the rather electronically similar interfaces of the films (Al2O3/Fe5GeTe2 and Fe5GeTe2/AlOx). Instead, the $D(t)$ data is more compatible with the constant behaviour of a bulk origin, with a $D=0.04$ to 0.05 mJ/m2, as measured in the thickest films. The weaker DMI in the 3-nm-thick film, which correlates with lower $M_{s}$ and $T_{c}$, can be attributed to weaker ferromagnetism in the ultrathin regime.
@@ 2 Discussion
Chakraborty et al. 15 attributed a possible bulk DMI in Fe3GeTe2 to intercalated Fe in the vdW gaps and Fe vacancies. In our films, however, intercalated Fe was not detected 20 (see Fig. S1 in S.I.); it would also raise the Curie temperature 14, which was not observed. Furthermore, Rutherford backscattering (RBS) analysis show a stoichiometric Fe content (Fe:Ge = 5.00±0.08) 20, indicating at most a very low concentration of Fe vacancies.
@@
A second class of mechanisms involves symmetry breaking within the Fe5GeTe2 vdW layers. This is supported by our RHEED observations of a $\sqrt{3}\times\sqrt{3}$ superstructure, which cannot derive from a specific vdW stacking. A non-centrosymmetric phase with a $\sqrt{3}\times\sqrt{3}$ superstructure was reported by Gopi et al. 11, who ascribed it to a pattern of 1/3 Te vacancies ordered along the perpendicular direction. Since this phase represented only a small fraction ($10^{-2}$) of the studied bulk crystals, it was deemed to have a negligible effect on DMI. While this phase might be more prevalent in epitaxial films, RBS showed no significant Te depletion in our films 20 (Te:Ge = $2.0\pm 0.08$ for the 12-nm-thick film), suggesting that this mechanism is unlikely to explain our results.
@@
Fe5GeTe2 is expected to have two partially-filled Fe sites (“split sites”) at different vertical positions within each monolayer, “D” and “U” (Fig. 1a). Several groups interpreted the $\sqrt{3}\times\sqrt{3}$ superstructure as the ordering of Fe split sites 13, 1. In the centrosymmetric structure, these sites are equally and randomly populated. However, ab initio calculations showed that the “UDD” and “UUD” configurations are energetically more favorable 33. While our RHEED measurements cannot determine whether this ordering extends beyond the top layer, electron microscopy analyses in Ref. 1 showed a similar short-range order in the bulk of Fe5GeTe2. Three groups have calculated the DMI assuming a perfectly ordered “UUU” arrangement in a monolayer, using different computational methods 34, 35, 5. Their predicted $|D|$ values vary significantly (0.76, 0.6, and 0.06 mJ/m2), despite all employing commonly used techniques. Our measured $|D|$ (0.04 – 0.05 mJ/m2) is comparable to the lowest theoretical estimate, and much lower than the highest prediction. This quantitative difference may be due to several reasons. First, the calculations do not account for temperature, which reduces the macroscopic $D$. Second, our films are not ordered in the “UUU” configuration since this would be inconsistent with the observed $\sqrt{3}\times\sqrt{3}$ superstructure. According to Ghosh et al. 35, the DMI of the “UUD” or “UDD” configurations is only slightly lower than that of the “UUU” configuration ($|D|=0.05$ instead of 0.06 mJ/m2), contrary to the naïve expectation of a 2/3 reduction. Finally, by symmetry, the “UUD” and “UDD” configurations contribute equally but with opposite sign DMI. While one of the two must be dominant to account for the observed DMI, the other may be present in a smaller amount and partially compensate the net DMI.
@@
Assuming that the DMI arises from an ordering of the Fe split sites, the sign of $D$ can be used to infer the dominant split-site configuration. Both refs. 34 and 5 found that a fully-ordered “UUU” configuration favours counterclockwise spinwaves, corresponding to a negative DMI. Therefore, the observed positive DMI suggests that the predominant ordering in our films is “UDD.”